The electron domain and molecular geometry of SO3 are ________. Question 38 options: A) trigonal bipyramidal, T-shaped B) octahedral, seesaw C) trigonal planar, trigonal planar D) trigonal bipyramidal, trigonal planar E) trigonal planar, bent
Answer #1

Angues SO3 S - 2,8,6 IG - 2,8 10: :s: :O: Ö=S -s= :0; Sterie No. of atomg bonded to centeal atom No. of lone pairs on centralElec+101) molecular Atomg bo- nded to centkal atomy Lone pains on central geomety geometry linear linear 2. O 2 lineaa linear